OpenStudy (anonymous):

Suppose cosA = 12/13 with 0º≤A≤90º. Suppose also that sinB = 8/17 with 90º≤B≤180º. Find cos(A - B).

OpenStudy (princeharryyy):

cos(A-B) = cosAcosB+sinAsinB

OpenStudy (princeharryyy):

solve using this @mashlyashela

OpenStudy (princeharryyy):

cosA = 12/13 sin^2A = 1 - cos^2A = 1 - 144/169 =25/169 so sin A = 5/13 ,A is acute(<90)

OpenStudy (princeharryyy):

similarly, cos^2B = 1 - sin^2B= 1 - 64/289 = 225/289 so cos B = - 15/17, since B is obtuse(if b would have been acute it would have been positive)

OpenStudy (princeharryyy):

Now cos ( A - B) = cos A cos B + sin A sin B = (12/13)*(- 15/17) + (5/13)*(8/17) = - 180/221 + 40/221 = - 140/221 ans

OpenStudy (princeharryyy):

@mashlyashela

OpenStudy (princeharryyy):

hope u understand @mashlyashela

OpenStudy (anonymous):

Yes thank you. Now I know what I did wrong

OpenStudy (princeharryyy):

where did u go wrong @mashlyashela

OpenStudy (anonymous):

I put 15/8 and 12/5

OpenStudy (princeharryyy):

from where did u got those @mashlyashela

OpenStudy (anonymous):

12/13 and 8/17, I put the fractions over the top numbers. 5/12, 15/8

OpenStudy (princeharryyy):

ohhhk !!! evry1 makes mistakes....

OpenStudy (princeharryyy):

@mashlyashela

OpenStudy (anonymous):

yup, lol thanks

OpenStudy (princeharryyy):

u r welcome....))) @mashlyashela